# Post-Training Language Models for Continual Relation Extraction

### v2

@@ 1 Introduction
To process unstructured non-stationary data and represent it in KGs, continuous learning techniques can identify entities and their relationships while handling real-time streaming data. In real-world scenarios, an RE model trained on stationary data might fail to recognize new relation types not introduced during its training phase. Streaming data, which is inherently non-stationary, requires continuous training and evaluation to detect these new relation types. Figure 1 illustrates how a model is incrementally trained on new RE tasks and evaluated on an incrementally expanding test set of relation types.
@@
In the rest of this paper, we first provide an overview about the preliminaries and related works in Section 2. We then introduce our methodology for CL of new relation types in Section 3. Following this, we evaluate our approach using two benchmark datasets and various LLMs, applying memory replay with incremental task learning, and compare it with previous methods in Section 4, and then conduct ablation study in Section 5. Following this, we discuss potential outcomes of our work with previous approach within the context of our research question in Section 6. Finally, we conclude by highlighting our findings and suggesting a future research direction in Section 7.
@@ 2 Related Works
Continual relation extraction (CRE) is a task that aims to train a model continuously on data containing new relation types, while preserving the knowledge of previously learned relations [26, 32] as illustrated in Figure 1. CRE task might be formalized by definitions based in [30]. Furthermore, an example of data for a relation type is defined as a tuple, $\mathbf{x=\langle\textit{sentence},\textit{head},\textit{tail}\rangle}$ where sentence is a textual sentence consisting of multiple tokens, head is the token(s) corresponding to the head entity, and tail is the token(s) corresponding to the tail entity. Additionally, Data $\textit{Data sequence}=\{X_{0},X_{1},\dots,X_{K}\}$ and Relation $\textit{Relation sequence}=\{R_{0},R_{1},\dots,R_{K}\}$ where $X_{K}$ contains the tuples for relations $R_{t}$ at time $t\leq K$, with $t$ indicating the time step. CRE approaches are commonly categorized into three types: (i) memory replay-based methods, (ii) dynamic architecture-based methods and (iii) regularization-based methods. In this section, we detail previously proposed methods designed to address the catastrophic forgetting phenomenon in CRE. The most commonly used benchmark datasets for CRE are TACRED [31] and FewRel [10].
@@ 3 Methodology
In this section, we present our methodology, integrating incremental task fine-tuning of pre-trained language models (PLMs) with memory replay techniques (see Section 3.1). The PLMs are fine-tuned using prompt instruction datasets derived from the original datasets through a structured prompt template selection process detailed in Section 3.2.
@@ 3.1 Continuous Learning with PLMs
The continuous fine-tuning process follows the steps outlined in Algorithm 1. As discussed in Section 1, memory replay (Lines 9–10 in Algorithm 1) is among the most effective strategies for mitigating catastrophic forgetting in continual learning. To select samples from the training data of previous tasks for memory replay after training on a new task (Lines 3–6), the most representative samples are identified by applying K-means clustering to the centroids of the clusters. This process utilizes embeddings from either the encoder or decoder of the trained model, depending on the architecture of the PLM. Specifically, decoder embeddings are used for decoder-only models, whereas encoder embeddings are applied for encoder-decoder models. After completing training on the new task, the validation dataset is used to optimize training parameters—such as the learning rate—based on validation loss (Lines 5–6) and this step is repeated throughout training (Lines 3-6). Validation datasets were not utilized for optimization in previous works [4, 32].
@@ 4       Retrieve current task $T_{k}$; /* $k$: position of current task */
6             Update $\theta$ with $\nabla L$ on ${D}^{k}_{train}$; /* $\nabla L$: gradient of classification loss on ${D}^{k}_{train}$ for $T_{k}$/
@@ 8            Adjust the learning rate based on $L_{valid}$;
13             Update $\theta$ with $\nabla L$ on $\widetilde{M}$; /* Fine-tune on memory $\widetilde{M}$/
@@ 16      $\widetilde{M}\leftarrow\widetilde{M}\cup M_{k}$; /* Add selected samples from $T_{k}$ to memory */
In reference to the details in Algorithm 1, the algorithm processes a stream of tasks and memory samples alongside the PLM model ($f_{\theta}$). The memory samples ($\widetilde{M}$) are initially empty and are dynamically selected from the training dataset (${D}^{k}_{train}$) to facilitate replay after training on the new task. The PLM model is incrementally trained on ${D}^{k}_{train}$ and validated on ${D}^{k}_{valid}$ per task as it arrives from the stream (Lines 3-6). Following the training of the PLM on the new task, memory samples are selected from the training data using K-means clustering (see Line 8). This selection leverages either the PLM’s encoder ($f_{\theta}^{encoder}$) or decoder to compute embeddings of the samples, previously trained in Lines 3–6. Before storing the selected samples, which represent the centroids of the clusters identified by K-means (Line 12), the memory samples from the previous task, denoted as $\widetilde{M}$, are replayed in Lines 9-10.
@@ 3.2 Prompt Template Selection
We investigate two prompt templates for incremental task fine-tuning without memory replay. The first template (see Figure 2(a)), derived from [21], is modified to incorporate relation types and employs conditional generation techniques [15]. In contrast, the second template are taken from [7], explicitly defines the task and specifies the head and tail entities, differentiating it from Template 1 (see Figure 2(a)). We then assess the model’s performance with both prompt templates, presenting the results in Section 4. We consider the best-performing template for subsequent experiments based on these results.
@@ 4 Evaluation
In this section, we first outline the experimental settings in Section 4.1 and then present the results in Section 4.2 and knowledge transfer analysis in Section 4.3, based on the performance metrics described in this section.
@@ 4.1 Experimental Settings
c.) Evaluation Metrics. The performance of our experiments are evaluated according to seen task accuracies across incremental task learning (ITL) as illustrated in Figure 1. Additionally, we also computed whole accuracy, average accuracy and backward knowledge transfer metrics. Whole Accuracy [31] is computed from the resulting model at the end of ITL on all test data of all tasks. Average Accuracy [31] is also computed from the resulting model trained on task k on all the test sets of all tasks seen up to stage k of ITL. Backward Knowledge Transfer [12] quantifies the degree of forgetting in previously learned tasks after learning a new one. This metric is crucial for determining whether backward knowledge transfer (bwt) occurs [22]. Notably, no prior work in continual relation extraction has reported on this metric [22], which is computed by $bwt=\frac{1}{N-1}\sum_{t=1}^{N}A_{N,t}-A_{t,t}$ where $A_{N,t}$ represents the test accuracy on the $t$-th task after sequential training on all $N$ tasks. All metrics results in this work are the mean of five runs, except for results shown in confusion matrices.
@@
e.) Prompt Template Selection. We compare two prompt templates for incremental task fine-tuning on TACRED across five runs. The results on the TACRED dataset indicate that for prompt type one, mean Whole Accuracy (w) is 92.7%, and the mean Average Accuracy (a) is 92.1%. In comparison, prompt type two achieves mean w of 90.5% and mean a of 90.6%. Therefore, we take into account prompt template one (see Figure 2(a)) to create prompt datasets from TACRED and FewRel benchmarks to fine-tune the aforementioned pre-trained language models.
@@ 4.2 Results
The ITL-based fine-tuned Flan T5 Base model achieves remarkable performance on TACRED, achieving impressive mean seen task accuracy with 95.8% on TACRED and with 89.58% on FewRel as shown in Table 1. It also demonstrates strong whole accuracy (w) and average accuracy (a), dealt with minimal forgetting on TACRED, as indicated by the mean backward knowledge transfer (bwt) of -0.2% in Table 2. Unfortunately, the model encounters significant forgetting challenges on FewRel, with a mean bwt of -1.75% (see Table 2). Furthermore, the mean w and a are 95.76% and 95.78% on TACRED, and 89.61% and 89.61% on FewRel, respectively as given in Table 2, indicating good performance on individual tasks with minimal catastrophic forgetting on TACRED.
In addition to Flan T5 Base, we evaluate the performance of Mistral on TACRED and FewRel as well. Mistral achieves positive bwt with 0.17% on TACRED in Table 2, indicating that it enhances the performance of previously learned tasks at the end of ITL. However, it encounters slight forgetting with a mean bwt of -1.0% on FewRel (see Table 2). Besides, the resulting models’ mean seen task accuracies on these benchmark datasets are 96.9% and 91.35%, respectively in Table 1 at the end of ITL where Mistral is used. The mean w and a are 96.89% and 96.76% on TACRED, and 94.93% and 94.93% on FewRel with this model in Table 1. Likewise, Mistral performs well on individual tasks with minimal catastrophic forgetting on TACRED when w and a metrics are considered.
Finally, we evaluate Llama2 in the context of ITL settings. In contrast to Flan T5 Base and Mistral, it does not achieve remarkable results on either dataset, with a mean seen task accuracy of 70.6% on TACRED and 71.29% on FewRel in Table 1 at the end of ITL. Interestingly, it demonstrates positive bwt on both datasets (see Table 2), a phenomenon rarely observed even in computer vision. Similar to its seen task accuracies, its mean w and a—71.17% and 70.86% on TACRED, and 71.29% for both metrics on FewRel—are lower than those of the other models (see Table 2), primarily due to hallucinating relation types by Llama2 like per:affiliate and per:columnist.
@@ 4.3 Knowledge Transfer Analysis
In this section, we analyze and visualize knowledge transfer in incremental task learning (ITL). To assess knowledge transfer across ITL, we examine the mean test accuracy of Task 1 from both FewRel and TACRED, utilizing three language models, as illustrated in Figure 3. The mean test accuracy of Task 1 shows a slight decline as Flan T5 Base progresses from training on Task 1 to Task 10, where TACRED is evaluated. Similarly, Mistral also experiences a bit forgetting on this dataset. In contrast, Llama2 achieves positive backward knowledge transfer over ITL, which enhances the model’s performance on earlier tasks (e.g., Task 1, as indicated in Figure 3(a)). Furthermore, we analyze the performance of these language models on the FewRel dataset as well. Flan T5 Base suffers from significant catastrophic forgetting throughout ITL, as depicted in Figure 3(b). We illustrate Flan T5 Base’s behavior with a confusion matrix in Figure 4, which highlights how test dataset of Task 1 performs during ITL. Flan T5 Base performs better on the Task 1’s test dataset after training on the Task 1’s training set (see Figure 4(a)); however, it generates hallucinated predictions—responses not among the predefined relation types—when encountering catastrophic forgetting, as shown in Figure 4(b). Note that the FewRel test dataset contains 140 samples per relation type. While calculating the number of predictions in the confusion matrices (see Figures 5 and 4), we exclude statistics on hallucinated predictions. Similar to the TACRED results, Mistral faces forgetting issues on FewRel, as illustrated in Figure 3(b). Llama2 demonstrates comparable positive backward knowledge transfer on FewRel (see Figure 3(b)), while reducing the hallucinated predictions by the completion of ITL. This reduction is evident from the decrease in false predictions throughout ITL, as shown in Figures 5(a) and 5(b). In conclusion, language models may generate the hallucinated predictions when they fail to transfer previously learned knowledge forward. Additionally, the number of hallucinated predictions tends to decrease when backward knowledge transfer occurs, as observed with Llama2.
@@ 5 Ablation Study
We examine how varying memory sample sizes (e.g., 5, 10, 15) affect incremental task fine-tuning using the Flan-T5 Base model on the TACRED dataset (see Table 3). We perform a two-tailed significance test with $H_{0}$: no difference between the condition without memory replay and memory size m, and $H_{1}$: a significant difference between m=0 (no replay) and m= 5, 10 or 15. The results show significant differences in memory sizes, with p-values of 0.0958 (m= 0 vs 5), 0.0509 (m= 0 vs 10), and 0.0262 (m= 0 vs 15), all below $\alpha$= 0.10 (level of significance), leading to the rejection of $H_{0}$.
@@ 6 Discussion
In this work, we investigate whether pre-trained language models influence knowledge transfer in CRE, seeking an answer to the research question: To what extent do pre-trained language models affect knowledge transfer, including backward transfer, in continual relation extraction? Positive backward knowledge transfer is observed with Llama2 on both datasets, reducing false predictions and hallucinations on earlier tasks in Section 4.3. In contrast, Flan T5 Base generates hallucinated predictions, particularly when encountering significant knowledge forgetting (see Figure 4). While custom CRE models produce false predictions among predefined relation types, language models tend to generate hallucinations beyond these predefined types, even when fine-tuned for a specific task.
Additionally, we also compare our findings with the state-of-the-art (SoTA) method [29]. Mistral outperforms Flan T5 Base and Llama2 on both datasets across evaluation metrics, although it exhibits slight forgetting on FewRel (see Table 2). Furthermore, it surpasses previously published SoTA results ([29]) on TACRED in terms of all evaluation metrics (see Tables 1 and 2) and exceeds KIP’s performance on FewRel in terms of seen task accuracy (see Table 1) at the end of ITL. However, Mistral ranks second when considering whole and average accuracies in Table 2. Notably, KIP leverages prompt-based approaches and multi-head scaled dot-product attention alongside memory replay.

### v3

@@ 1 Introduction
To process unstructured non-stationary data and represent it in KGs, continuous learning techniques can identify entities and their relationships while handling real-time streaming data. In real-world scenarios, an RE model trained on stationary data might fail to recognize new relation types not introduced during its training phase. Streaming data, which is inherently non-stationary, requires continuous training and evaluation to detect these new relation types. Figure˜1 illustrates how a model is incrementally trained on new RE tasks and evaluated on an incrementally expanding test set of relation types.
@@
In the rest of this paper, we first provide an overview about the preliminaries and related works in Section˜2. We then introduce our methodology for CL of new relation types in Section˜3. Following this, we evaluate our approach using two benchmark datasets and various LLMs, applying memory replay with incremental task learning, and compare it with previous methods in Section˜4, and then conduct ablation study in Section˜5. Following this, we discuss potential outcomes of our work with previous approach within the context of our research question in Section˜6. Finally, we conclude by highlighting our findings and suggesting a future research direction in Section˜7.
@@ 2 Related Works
Continual relation extraction (CRE) is a task that aims to train a model continuously on data containing new relation types, while preserving the knowledge of previously learned relations [26, 32] as illustrated in Figure˜1. CRE task might be formalized by definitions based in [30]. Furthermore, an example of data for a relation type is defined as a tuple, $\mathbf{x=\langle\textit{sentence},\textit{head},\textit{tail}\rangle}$ where sentence is a textual sentence consisting of multiple tokens, head is the token(s) corresponding to the head entity, and tail is the token(s) corresponding to the tail entity. Additionally, $\textit{Data sequence}=\{X_{0},X_{1},\dots,X_{K}\}$ and $\textit{Relation sequence}=\{R_{0},R_{1},\dots,R_{K}\}$ where $X_{K}$ contains the tuples for relations $R_{t}$ at time $t\leq K$, with $t$ indicating the time step. CRE approaches are commonly categorized into three types: (i) memory replay-based methods, (ii) dynamic architecture-based methods and (iii) regularization-based methods. In this section, we detail previously proposed methods designed to address the catastrophic forgetting phenomenon in CRE. The most commonly used benchmark datasets for CRE are TACRED [31] and FewRel [10].
@@ 3 Methodology
In this section, we present our methodology, integrating incremental task fine-tuning of pre-trained language models (PLMs) with memory replay techniques (see Section˜3.1). The PLMs are fine-tuned using prompt instruction datasets derived from the original datasets through a structured prompt template selection process detailed in Section˜3.2.
@@ 3.1 Continuous Learning with PLMs
The continuous fine-tuning process follows the steps outlined in Algorithm˜1. As discussed in Section˜1, memory replay (Lines 9–10 in Algorithm˜1) is among the most effective strategies for mitigating catastrophic forgetting in continual learning. To select samples from the training data of previous tasks for memory replay after training on a new task (Lines 3–6), the most representative samples are identified by applying K-means clustering to the centroids of the clusters. This process utilizes embeddings from either the encoder or decoder of the trained model, depending on the architecture of the PLM. Specifically, decoder embeddings are used for decoder-only models, whereas encoder embeddings are applied for encoder-decoder models. After completing training on the new task, the validation dataset is used to optimize training parameters—such as the learning rate—based on validation loss (Lines 5–6) and this step is repeated throughout training (Lines 3-6). Validation datasets were not utilized for optimization in previous works [4, 32].
@@ 4   Retrieve current task $T_{k}$; /* $k$: position of current task */
6      Update $\theta$ with $\nabla L$ on ${D}^{k}_{train}$; /* $\nabla L$: gradient of classification loss on ${D}^{k}_{train}$ for $T_{k}$ */
@@ 8     Adjust the learning rate based on $L_{valid}$;
13      Update $\theta$ with $\nabla L$ on $\widetilde{M}$; /* Fine-tune on memory $\widetilde{M} *$/
@@ 16 $\widetilde{M}\leftarrow\widetilde{M}\cup M_{k}$; /* Add selected samples from $T_{k}$ to memory */
In reference to the details in Algorithm˜1, the algorithm processes a stream of tasks and memory samples alongside the PLM model ($f_{\theta}$). The memory samples ($\widetilde{M}$) are initially empty and are dynamically selected from the training dataset (${D}^{k}_{train}$) to facilitate replay after training on the new task. The PLM model is incrementally trained on ${D}^{k}_{train}$ and validated on ${D}^{k}_{valid}$ per task as it arrives from the stream (Lines 3-6). Following the training of the PLM on the new task, memory samples are selected from the training data using K-means clustering (see Line 8). This selection leverages either the PLM’s encoder ($f_{\theta}^{encoder}$) or decoder to compute embeddings of the samples, previously trained in Lines 3–6. Before storing the selected samples, which represent the centroids of the clusters identified by K-means (Line 12), the memory samples from the previous task, denoted as $\widetilde{M}$, are replayed in Lines 9-10.
@@ 3.2 Prompt Template Selection
We investigate two prompt templates for incremental task fine-tuning without memory replay. The first template (see Figure˜2(a)), derived from [21], is modified to incorporate relation types and employs conditional generation techniques [15]. In contrast, the second template are taken from [7], explicitly defines the task and specifies the head and tail entities, differentiating it from Template 1 (see Figure˜2(a)). We then assess the model’s performance with both prompt templates, presenting the results in Section˜4. We consider the best-performing template for subsequent experiments based on these results.
@@ 4 Evaluation
In this section, we first outline the experimental settings in Section˜4.1 and then present the results in Section˜4.2 and knowledge transfer analysis in Section˜4.3, based on the performance metrics described in this section.
@@ 4.1 Experimental Settings
c.) Evaluation Metrics. The performance of our experiments are evaluated according to seen task accuracies across incremental task learning (ITL) as illustrated in Figure˜1. Additionally, we also computed whole accuracy, average accuracy and backward knowledge transfer metrics. Whole Accuracy [31] is computed from the resulting model at the end of ITL on all test data of all tasks. Average Accuracy [31] is also computed from the resulting model trained on task k on all the test sets of all tasks seen up to stage k of ITL. Backward Knowledge Transfer [12] quantifies the degree of forgetting in previously learned tasks after learning a new one. This metric is crucial for determining whether backward knowledge transfer (bwt) occurs [22]. Notably, no prior work in continual relation extraction has reported on this metric [22], which is computed by $bwt=\frac{1}{N-1}\sum_{t=1}^{N}A_{N,t}-A_{t,t}$ where $A_{N,t}$ represents the test accuracy on the $t$-th task after sequential training on all $N$ tasks. All metrics results in this work are the mean of five runs, except for results shown in confusion matrices.
@@
e.) Prompt Template Selection. We compare two prompt templates for incremental task fine-tuning on TACRED across five runs. The results on the TACRED dataset indicate that for prompt type one, mean Whole Accuracy (w) is 92.7%, and the mean Average Accuracy (a) is 92.1%. In comparison, prompt type two achieves mean w of 90.5% and mean a of 90.6%. Therefore, we take into account prompt template one (see Figure˜2(a)) to create prompt datasets from TACRED and FewRel benchmarks to fine-tune the aforementioned pre-trained language models.
@@ 4.2 Results
The ITL-based fine-tuned Flan T5 Base model achieves remarkable performance on TACRED, achieving impressive mean seen task accuracy with 95.8% on TACRED and with 89.58% on FewRel as shown in Table˜1. It also demonstrates strong whole accuracy (w) and average accuracy (a), dealt with minimal forgetting on TACRED, as indicated by the mean backward knowledge transfer (bwt) of -0.2% in Table˜2. Unfortunately, the model encounters significant forgetting challenges on FewRel, with a mean bwt of -1.75% (see Table˜2). Furthermore, the mean w and a are 95.76% and 95.78% on TACRED, and 89.61% and 89.61% on FewRel, respectively as given in Table˜2, indicating good performance on individual tasks with minimal catastrophic forgetting on TACRED.
In addition to Flan T5 Base, we evaluate the performance of Mistral on TACRED and FewRel as well. Mistral achieves positive bwt with 0.17% on TACRED in Table˜2, indicating that it enhances the performance of previously learned tasks at the end of ITL. However, it encounters slight forgetting with a mean bwt of -1.0% on FewRel (see Table˜2). Besides, the resulting models’ mean seen task accuracies on these benchmark datasets are 96.9% and 91.35%, respectively in Table˜1 at the end of ITL where Mistral is used. The mean w and a are 96.89% and 96.76% on TACRED, and 94.93% and 94.93% on FewRel with this model in Table˜1. Likewise, Mistral performs well on individual tasks with minimal catastrophic forgetting on TACRED when w and a metrics are considered.
Finally, we evaluate Llama2 in the context of ITL settings. In contrast to Flan T5 Base and Mistral, it does not achieve remarkable results on either dataset, with a mean seen task accuracy of 70.6% on TACRED and 71.29% on FewRel in Table˜1 at the end of ITL. Interestingly, it demonstrates positive bwt on both datasets (see Table˜2), a phenomenon rarely observed even in computer vision. Similar to its seen task accuracies, its mean w and a—71.17% and 70.86% on TACRED, and 71.29% for both metrics on FewRel—are lower than those of the other models (see Table˜2), primarily due to hallucinating relation types by Llama2 like per:affiliate and per:columnist.
@@ 4.3 Knowledge Transfer Analysis
In this section, we analyze and visualize knowledge transfer in incremental task learning (ITL). To assess knowledge transfer across ITL, we examine the mean test accuracy of Task 1 from both FewRel and TACRED, utilizing three language models, as illustrated in Figure˜3. The mean test accuracy of Task 1 shows a slight decline as Flan T5 Base progresses from training on Task 1 to Task 10, where TACRED is evaluated. Similarly, Mistral also experiences a bit forgetting on this dataset. In contrast, Llama2 achieves positive backward knowledge transfer over ITL, which enhances the model’s performance on earlier tasks (e.g., Task 1, as indicated in Figure˜3(a)). Furthermore, we analyze the performance of these language models on the FewRel dataset as well. Flan T5 Base suffers from significant catastrophic forgetting throughout ITL, as depicted in Figure˜3(b). We illustrate Flan T5 Base’s behavior with a confusion matrix in Figure˜4, which highlights how test dataset of Task 1 performs during ITL. Flan T5 Base performs better on the Task 1’s test dataset after training on the Task 1’s training set (see Figure˜4(a)); however, it generates hallucinated predictions—responses not among the predefined relation types—when encountering catastrophic forgetting, as shown in Figure˜4(b). Note that the FewRel test dataset contains 140 samples per relation type. While calculating the number of predictions in the confusion matrices (see Figures˜5 and 4), we exclude statistics on hallucinated predictions. Similar to the TACRED results, Mistral faces forgetting issues on FewRel, as illustrated in Figure˜3(b). Llama2 demonstrates comparable positive backward knowledge transfer on FewRel (see Figure˜3(b)), while reducing the hallucinated predictions by the completion of ITL. This reduction is evident from the decrease in false predictions throughout ITL, as shown in Figures˜5(a) and 5(b). In conclusion, language models may generate the hallucinated predictions when they fail to transfer previously learned knowledge forward. Additionally, the number of hallucinated predictions tends to decrease when backward knowledge transfer occurs, as observed with Llama2.
@@ 5 Ablation Study
We examine how varying memory sample sizes (e.g., 5, 10, 15) affect incremental task fine-tuning using the Flan-T5 Base model on the TACRED dataset (see Table˜3). We perform a two-tailed significance test with $H_{0}$: no difference between the condition without memory replay and memory size m, and $H_{1}$: a significant difference between m=0 (no replay) and m= 5, 10 or 15. The results show significant differences in memory sizes, with p-values of 0.0958 (m= 0 vs 5), 0.0509 (m= 0 vs 10), and 0.0262 (m= 0 vs 15), all below $\alpha$= 0.10 (level of significance), leading to the rejection of $H_{0}$.
@@ 6 Discussion
In this work, we investigate whether pre-trained language models influence knowledge transfer in CRE, seeking an answer to the research question: To what extent do pre-trained language models affect knowledge transfer, including backward transfer, in continual relation extraction? Positive backward knowledge transfer is observed with Llama2 on both datasets, reducing false predictions and hallucinations on earlier tasks in Section˜4.3. In contrast, Flan T5 Base generates hallucinated predictions, particularly when encountering significant knowledge forgetting (see Figure˜4). While custom CRE models produce false predictions among predefined relation types, language models tend to generate hallucinations beyond these predefined types, even when fine-tuned for a specific task.
Additionally, we also compare our findings with the state-of-the-art (SoTA) method [29]. Mistral outperforms Flan T5 Base and Llama2 on both datasets across evaluation metrics, although it exhibits slight forgetting on FewRel (see Table˜2). Furthermore, it surpasses previously published SoTA results ([29]) on TACRED in terms of all evaluation metrics (see Tables˜1 and 2) and exceeds KIP’s performance on FewRel in terms of seen task accuracy (see Table˜1) at the end of ITL. However, Mistral ranks second when considering whole and average accuracies in Table˜2. Notably, KIP leverages prompt-based approaches and multi-head scaled dot-product attention alongside memory replay.